# Leading by the Nodes: A Survey of Film Industry Network Analysis and Datasets

### v1

@@ 5 Discussion and Conclusions
Second, we note that a disproportionate amount of the literature (especially the formative literature for theory-building in this area) is focused on the film industry of the United States, and more work is needed to see whether other film industries are governed by the same network principles. This focus on Hollywood may well be related to the geographic profile of who is contributing to this emerging literature, as our analysis found that the co-authorship network is dominated by scholars from the Minority World (particularly the United States, Europe and Australia). In this respect, it is encouraging to see that in recent years there has been a growth in the number of contributions to this research area from authors in the Majority World, including analyses of film industries outside of the dominant global export markets. However, our citation network analysis shows that recognition and scholarly engagement are uneven, with authors in the Minority World citing the contributions of scholars in the Majority World far less than they are cited by them (Fig. 5a and Fig. 5b).
@@
Fourth, our review highlights that, with a few valuable exceptions, very little scholarship has been done which looks at inequalities in the film industry from a relational perspective. Qualitative and non-relational quantitative evidence (not reviewed in this paper) persistently highlights how inequitable these industries and their workforces are. Moreover, the foundational work we reviewed by Faulkner and Jones provides rich evidence of how important reputational and relational mechanisms are for structuring opportunities in the project-based cultural industries. These factors intersect with known network mechanisms of exclusion such as homophily and familiarity to systemically disadvantage some people and benefit others. Despite this, relatively few papers take a network analytic approach to studying film industry inequalities, which makes it harder for policies to be developed which aim to redress the network-based behaviours which shape the industry.
@@
Finally, more engagement with existing scholarship is needed, especially among researchers from computational disciplines who wish to use the film industry as a case study for advanced methods. There is enough literature that we should expect more cross-citation within the network. However, papers continue to appear which apply network analysis to the film industry yet do not acknowledge any of the earlier work in this area. For the research area to continue to develop, scholars need to properly recognize and engage with the vast amount of work already done. We hope that this survey paper can help simplify this process for scholars wishing to contribute further to this fast-growing topic.
Together, these findings contribute a synthesis of an important and popular area of network science application that has not yet been systematically reviewed to find coherent themes and gaps. We have highlighted several ways in which applied network scientists can advance understanding of the film industry, and flagged areas that our survey reveals as being in particular need of more attention. There are also several limitations to the present survey, and in particular to our ability to analyze the networks derived from the set of papers we reviewed. Firstly, the restrictions we placed on our literature search, including the decision to only review publications in English, helped ensure that a survey was feasible and bounded, but it inevitably means that we will not have included some valuable contributions in this space. Furthermore, co-authorship data, while informative, constitutes just one of several indicators of scientific collaboration. Not all collaborative endeavors culminate in publications, and the co-authorship or citation of papers does not invariably imply knowledge-sharing among authors.

### v2

@@ 4 Discussion and Conclusions
Second, we note that a disproportionate amount of the literature (especially the formative literature for theory-building in this area) is focused on the film industry of the United States, and more work is needed to see whether other film industries are governed by the same network principles. This focus on Hollywood may well be related to the geographic profile of who is contributing to this emerging literature, as a review of the institutional affiliations of authors shows that this literature is dominated by scholars from the Minority World (particularly the United States, Europe and Australia). In this respect, it is encouraging to see that in recent years there has been a growth in the number of contributions to this research area from authors in the Majority World, including analyses of film industries outside of the dominant global export markets. However, we note that the flow of recognition and scholarly engagement is uneven, with authors in the Minority World citing the contributions of scholars in the Majority World far less than they are cited by them.
@@
Fourth, our review highlights that, with a few valuable exceptions, very little scholarship has been done which looks at inequalities in the film industry from a relational perspective. Qualitative and non-relational quantitative evidence (not reviewed in this paper) persistently highlights how inequitable these industries and their workforces are. Moreover, the foundational work we reviewed by Faulkner and Jones provides rich evidence of how important reputational and relational mechanisms are for structuring opportunities in project-based cultural industries. These factors intersect with known network mechanisms of exclusion such as homophily and familiarity to systemically disadvantage some people and benefit others. Despite this, relatively few papers take a network analytic approach to studying film industry inequalities, which in turn makes it harder to develop evidence-based policies that aim to redress the industry’s uneven relationships.
@@
Finally, more engagement with existing scholarship is needed, especially among researchers from computational disciplines who wish to use the film industry as a case study for advanced methods. There is enough literature that we should expect more cross-citation among the contributions. However, papers continue to appear which apply network analysis to the film industry and yet do not acknowledge any of the earlier work in this area. For the research area to continue to develop, scholars need to properly recognize, engage with and build on the vast amount of scholarship already undertaken. We hope that this survey paper can help simplify this process for scholars wishing to contribute further to this fast-growing topic.
Together, these findings contribute a synthesis of an important and popular area of network science application that has not yet been systematically reviewed to find coherent themes and gaps. We have highlighted several ways in which applied network scientists can advance understanding of the film industry, and flagged areas that our survey reveals as being in particular need of more attention. There are also several limitations to the present survey. Firstly, the restrictions we placed on our literature search, including the decision to only review publications in English, helped ensure that a survey was feasible and bounded, but it inevitably means that we will not have included some valuable contributions in this space. Furthermore, as the global film industry grapples with technological, business and social challenges we acknowledge the types of nodes we have outlined and their relational dimensions will also evolve and we look forward to the terms of this survey being extended in future iterations of film industry network analysis.